# Asymptotics of spherical dynamos exhibiting a small-scale MAC balance

### v1

@@ 5 Conclusion
The kinetic energy spectra computed from the non-axisymmetric motions show that the simulations are characterised by a broad range of length scales in which different regions of the spectra exhibit different asymptotic dependencies on the Ekman number. Similar behaviour was observed in non-magnetic spherical convection with stress-free boundary conditions (Nicoski et al., 2024). The largest length scales, as computed from the peak of the kinetic energy spectra, show an approximate scaling of $O(Ek^{1/6})$, whereas scales smaller than this are well characterised by the $O(Ek^{1/3})$ scaling. The flow speeds, in agreement with QG theory, follow a $Re=O\left(Ek^{-1/3}\right)$ scaling. Prior work has suggested that the presence of $O(Ek^{1/3})$ length scales in simulations is an indication that viscosity plays a dominant role in the dynamics (King & Buffett, 2013). We stress that this conclusion is not generally true. Rather, this scaling arises from the asymptotic nature of many different terms in the governing equations and does not in itself indicate a dominant role of viscosity. Indeed, as shown here, a length scale derived from the advection term also exhibits this same asymptotic scaling.
@@
Length scales computed from the magnetic field are approximately characterised by a $O(Ek^{1/6})$ scaling, i.e. substantially weaker than the dominant $O(Ek^{1/3})$ scaling observed in the velocity field. As discussed in prior work, the magnetic induction equation can be used to deduce a ohmic dissipation length scale that follows a $Rm^{-1/2}$ scaling (e.g. Christensen & Tilgner, 2004). Here we have shown that this scaling is equivalent to an asymptotic dependence of $O(Ek^{1/6})$ for $Pm=O(1)$. Further work is necessary to determine the reason for the differences in length scales in the magnetic and velocity fields.
@@
Similar to other studies (e.g. Dormy, 2016; Menu et al., 2020), we find that increasing the magnetic Prandtl number can greatly increase the Lorentz force, and we also find that the buoyancy force can increase with $Pm$. This seems to be in contradiction to diffusion-free scaling arguments, which assume that the force balance is independent of $Pm$ (e.g. Davidson, 2013). The question remains as to whether dynamos are independent of magnetic diffusivity at more extreme parameter values, which could be tested directly by varying $Pm$ for low Ekman number high Rayleigh number simulations. If so, it would be interesting to find at what parameter values simulations become independent of $Pm$, which could help guide future investigations.
@@
It is often assumed that advection plays no role in the geodynamo because the Rossby number in the outer core is small. However, this argument is oversimplified given what is known about the perturbative dynamics of rapidly rotating convection; while both inertia and advection are small relative to the leading order geostrophic balance, they nevertheless play crucial roles in the dynamics of geostrophic convective turbulence, even for flows with an asymptotically small Rossby number (e.g. Julien et al., 2012). Moreover, a plethora of waves and time-dependent flows are thought to be important in the core (e.g. Finlay et al., 2023). The simulations analysed here confirm that advection is small relative to the leading order forces, though we find that it is comparable in magnitude to, and shows the same asymptotic dependence as, both inertia and the viscous force. This finding suggests that the temporal dynamics are characterised by an $O(Ek^{2/3})$ timescale (in our non-dimensional units), which is equivalent to the convective Rossby wave timescale from asymptotic theory (Roberts, 1968), and found to be the dominant timescale in prior studies of turbulent rotating convection (Oliver et al., 2025). This result is surprising in light of the observed force balance (MAC). On the other hand, the prevalence of the $Ek^{1/3}$ in the velocity field essentially requires this to be the dominant inertial timescale.
@@
Aubert et al. (2017) and related subsequent work (e.g. Schwaiger et al., 2019, 2021) utilised spherical harmonic power spectra of the forces in the governing equations in an effort to better understand the relationship between force balances and length scales in both non-magnetic rotating convection and rotating convection-driven dynamos. By basing their analysis on the spectral representation of the forces, as opposed to physical space, they conclude that dynamos with sufficiently strong magnetic field are geostrophically balanced to leading order on large length scales (small spherical harmonic degrees), and the Lorentz force can act to perturb this balance at some intermediate length scale. The authors refer to such dynamos as ‘QG-MAC’, even though the physical space force balance can be MAC. One of the reasons for the discrepancy between how our force balances are characterised in physical space and their corresponding spectral analysis is due to the behaviour of the force spectra, which are rarely characterised by well-defined peaks, implying that a large number of spherical harmonic degrees contribute to the dynamics. Indeed, it is the sum over all of these degrees that leads to the correct magnitude of the forces, and therefore the force balance, in physical space. It is also important to note that the axisymmetric component of the forces is not removed from their spectra, even though the axisymmetric and non-axisymmetric dynamics can be characterised by distinct force balances, as shown in Calkins et al. (2021) and the present work. The axisymmetric component of the forces tends to be dominated by even spherical harmonic degrees and can have a strong influence on the spectra both in terms of their amplitude and shape, as indicated by the ‘sawtooth’ pattern in their spectra. Furthermore, Teed & Dormy (2023) showed that while the spherical harmonic representation of the forces may show distinct ‘crossings’ in spectral space, which are used as proxies for length scales, the spherical harmonic representation of the curl of the forces (i.e. using the vorticity equation) lacks such spectral crossings, which represents a possible issue given that dynamo scaling laws often rely on assuming balances in the vorticity equation.
It is interesting to note one implication of the scaling that was found for the advection and viscous dissipation length scales. For the Earth, the Rossby number is approximately $Ro=10^{-6}$. We can define the small-scale Rossby number as $Ro_{\ell}=u/\ell\Omega$, where $\ell$ is the small-scale length scale relevant to the advective term of the momentum equation. If we assume the scaling $\ell/H\sim Ek^{1/3}$ we found for our dynamo simulations holds for Earth-like values and does not vary much with $\widetilde{Ra}$, then the small-scale Rossby number of the Earth would be approximately $Ro_{\ell}=Ro(\ell/H)^{-1}\sim 10^{-6}(10^{-15})^{-1/3}$, which implies $Ro_{\ell}\sim 0.1$. This is similar to estimates made in previous work using different arguments (e.g. Olson & Christensen, 2006). This value is near the cutoff between dipolar and non-dipolar dynamos that Christensen & Aubert (2006) found at $Ro_{\ell}\sim 0.12$, so Olson & Christensen (2006) pointed out that the Earth may be near the transition to the multipolar state, which could be why the Earth’s magnetic field reverses. However, some recent studies have pointed out that near the $Ro_{\ell}\sim 0.12$ transition where reversals occur, the magnetic to kinetic energy in dynamo simulations is approximately less than one (Zaire et al., 2022; Tassin et al., 2021). Since the Earth is believed to have much greater magnetic than kinetic energy in the outer core (Roberts & King, 2013), this sheds some doubt on the Earth being near the dipolar multipolar transition. On the other hand, if the length scale associated with the Lorentz force and the length scale associated with the advection term in the momentum equation are significantly different for the Earth, as suggested by our results, then the magnetic to kinetic energy could be much greater than one even with $Ro_{\ell}\sim 0.12$. This could be the case if the length scale of the Lorentz force is approximately $Rm^{-1/2}$ and the length scale of the advection term in the momentum equation is roughly $Ek^{1/3}$, for instance. Simulations carried out at smaller Ekman numbers would help test this hypothesis.
@@ Acknowledgements
This work was supported by the Geophysics Program at the National Science Foundation (NSF) through grant numbers EAR-1945270 and EAR-2201595. CD acknowledges support from the Natural Environment Research Council of the UK under grant number NE/V010867/1. Simulations were carried out on the Alpine supercomputer at the University of Colorado Boulder, the Anvil supercomputer at Purdue University and Stampede2 at the Texas Advanced Computing Center (TACC). Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260. Anvil and Stampede2 were made available through allocation PHY180013 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by NSF grants 2138259, 2138286, 2138307, 2137603 and 2138296. Visualisations were made using paraview (Ahrens et al., 2005).
DATA AVAILABILITY
The data used in this study will be shared upon reasonable request to the corresponding author.

### v3

@@ 5 Conclusion
The kinetic energy spectra computed from the non-axisymmetric motions show that the simulations are characterised by a broad range of length scales in which different regions of the spectra exhibit different asymptotic dependencies on the Ekman number. Similar behaviour was observed in non-magnetic spherical convection with stress-free boundary conditions [jN24]. The largest length scales, as computed from the peak of the kinetic energy spectra, show an approximate scaling of $O(Ek^{1/6})$, whereas scales smaller than this are well characterised by the $O(Ek^{1/3})$ scaling. The flow speeds, in agreement with QG theory, follow a $Re=O\left(Ek^{-1/3}\right)$ scaling. Prior work has suggested that the presence of $O(Ek^{1/3})$ length scales in simulations is an indication that viscosity plays a dominant role in the dynamics [eK13b]. We stress that this conclusion is not generally true. Rather, this scaling arises from the asymptotic nature of many different terms in the governing equations and does not in itself indicate a dominant role of viscosity. Indeed, as shown here, a length scale derived from the advection term also exhibits this same asymptotic scaling.
@@
Length scales computed from the magnetic field are approximately characterised by a $O(Ek^{1/6})$ scaling, i.e. substantially weaker than the dominant $O(Ek^{1/3})$ scaling observed in the velocity field. As discussed in prior work, the magnetic induction equation can be used to deduce a ohmic dissipation length scale that follows a $Rm^{-1/2}$ scaling [uC04, e.g.]. Here we have shown that this scaling is equivalent to an asymptotic dependence of $O(Ek^{1/6})$ for $Pm=O(1)$. Further work is necessary to determine the reason for the differences in length scales in the magnetic and velocity fields.
@@
Similar to other studies [eD16, mM20, e.g.], we find that increasing the magnetic Prandtl number can greatly increase the Lorentz force, and we also find that the buoyancy force can increase with $Pm$. This seems to be in contradiction to diffusion-free scaling arguments, which assume that the force balance is independent of $Pm$ [pD13b, e.g.]. The question remains as to whether dynamos are independent of magnetic diffusivity at more extreme parameter values, which could be tested directly by varying $Pm$ for low Ekman number high Rayleigh number simulations. If so, it would be interesting to find at what parameter values simulations become independent of $Pm$, which could help guide future investigations.
@@
It is often assumed that advection plays no role in the geodynamo because the Rossby number in the outer core is small. However, this argument is oversimplified given what is known about the perturbative dynamics of rapidly rotating convection; while both inertia and advection are small relative to the leading order geostrophic balance, they nevertheless play crucial roles in the dynamics of geostrophic convective turbulence, even for flows with an asymptotically small Rossby number [kJ12, e.g.]. Moreover, a plethora of waves and time-dependent flows are thought to be important in the core [cF23, e.g.]. The simulations analysed here confirm that advection is small relative to the leading order forces, though we find that it is comparable in magnitude to, and shows the same asymptotic dependence as, both inertia and the viscous force. This finding suggests that the temporal dynamics are characterised by an $O(Ek^{2/3})$ timescale (in our non-dimensional units), which is equivalent to the convective Rossby wave timescale from asymptotic theory [pR68], and found to be the dominant timescale in prior studies of turbulent rotating convection [tO25]. This result is surprising in light of the observed force balance (MAC). On the other hand, the prevalence of the $Ek^{1/3}$ in the velocity field essentially requires this to be the dominant inertial timescale.
@@
[jA17] and related subsequent work [tS19, tS21, e.g.] utilised spherical harmonic power spectra of the forces in the governing equations in an effort to better understand the relationship between force balances and length scales in both non-magnetic rotating convection and rotating convection-driven dynamos. By basing their analysis on the spectral representation of the forces, as opposed to physical space, they conclude that dynamos with sufficiently strong magnetic field are geostrophically balanced to leading order on large length scales (small spherical harmonic degrees), and the Lorentz force can act to perturb this balance at some intermediate length scale. The authors refer to such dynamos as ‘QG-MAC’, even though the physical space force balance can be MAC. One of the reasons for the discrepancy between how our force balances are characterised in physical space and their corresponding spectral analysis is due to the behaviour of the force spectra, which are rarely characterised by well-defined peaks, implying that a large number of spherical harmonic degrees contribute to the dynamics. Indeed, it is the sum over all of these degrees that leads to the correct magnitude of the forces, and therefore the force balance, in physical space. It is also important to note that the axisymmetric component of the forces is not removed from their spectra, even though the axisymmetric and non-axisymmetric dynamics can be characterised by distinct force balances, as shown in [mC21] and the present work. The axisymmetric component of the forces tends to be dominated by even spherical harmonic degrees and can have a strong influence on the spectra both in terms of their amplitude and shape, as indicated by the ‘sawtooth’ pattern in their spectra. Furthermore, [rT23] showed that while the spherical harmonic representation of the forces may show distinct ‘crossings’ in spectral space, which are used as proxies for length scales, the spherical harmonic representation of the curl of the forces (i.e. using the vorticity equation) lacks such spectral crossings, which represents a possible issue given that dynamo scaling laws often rely on assuming balances in the vorticity equation.
@@ Acknowledgements
This work was supported by the Geophysics Program at the National Science Foundation (NSF) through grant numbers EAR-1945270 and EAR-2201595. CD acknowledges support from the Natural Environment Research Council of the UK under grant number NE/V010867/1. Simulations were carried out on the Alpine supercomputer at the University of Colorado Boulder, the Anvil supercomputer at Purdue University and Stampede2 at the Texas Advanced Computing Center (TACC). Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260. Anvil and Stampede2 were made available through allocation PHY180013 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by NSF grants 2138259, 2138286, 2138307, 2137603 and 2138296. Visualisations were made using paraview [jA05b].